OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

how do i find the equation?

OpenStudy (trojanpoem):

Give me a sec , I misread the question. sorry.

OpenStudy (trojanpoem):

First , get the slope from the equation , I wrote before.

OpenStudy (trojanpoem):

\[m = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (trojanpoem):

You want direct answer or to learn ?

OpenStudy (anonymous):

direct answer please

OpenStudy (trojanpoem):

m = 1-2/5+2 = -1/7 y -1 /x -5 = -1/7 so -x = 7y - 12 x = -7y +12

OpenStudy (trojanpoem):

To vertify if it's right , just put on it the other point E -2 = -14 + 12 = -2 ( we are right mostly ).